On another planet, a marble is released from rest at the top of a high cliff. It falls 4.00 m in the first 1 s of its motion. Through what additional distance does it fall in the next 1 s?
1
Expert's answer
2020-10-07T07:25:40-0400

Distance traveled in first second "=s=ut+ \frac{1}{2}at^2=\frac{1}{2}at^2" "(" As "u=0)"

As "s_1= 4 \ m" for "t=1\ s"

So, "\frac{1}{2}a\times 1^2= 4\implies a= 8 \ m/s^2"

Now ,to find the distance traveled in first two seconds.

"t= 2 \ s, a= 8\ m/s^2"

"s_2= \frac{1}{2}at^2= \frac{1}{2}\times 8\times 2^2= 16\ s"

So distance traveled in "2"nd second is the difference between distance traveled by particle in first two seconds and first second.

Required distance"= s_2-s_1= 16-4=12\ m"